OpenStudy (anonymous):

How would I start working a first order differential equation (with an initial value) like this? (x^3 - y) dx + x dy = 0, y(1) = 3

OpenStudy (anonymous):

The given equation is : \[x ^{3} + x*dy(x)/dx - y(x) = 0\] = \[dy(x)/dx - y(x)/x = -x ^{2}\] Now put, \[\mu(x) = e ^{\int\limits -1/x dx} = 1/x\] Multiply both sides of the equation by \[\mu(x)\]to get : \[[dy(x)/dx]/x - y(x)/x ^{2} = -x\] = \[[dy(x)/dx]/x +d/dx(1/x)y(x) = -x\] Now apply the reverse product rule to get: \[d/dx(y(x)/x) = -x\] or,\[\int\limits d/dx(y(x)/x) dx = \int\limits -x dx\] or,\[y(x)/x = -x ^{2}/2 +c _{1}\] Now you can proceed easily!